Calculate cost of goods sold and ending inventory for Emergicare’s bandages orders using FIFO, LIFO and average cost. There are 34 units in ending inventory. (Do not round intermediate calculations, such as average cost per unit values. Round your answers to the nearest cent.)

Date Units purchased Cost
per unit
Total cost
January 1 51 $ 7.10 $ 362.10
April 1 46 6.35 292.10
June 1 61 6.10 372.10
September 1 56 6.60 369.60
Total 214 $ 1,395.90
Cost of goods sold Ending inventory
FIFO $ $
LIFO $ $
Average cost $ $
Answer #1

The answer has been presented in the supporting sheet. All the parts has been solved with detailed explanation and format. For detailed answer refer to the supporting sheet.

Answer 3 FIFO 4 Cost of ending inventory 5 = No. of units in ending inventory* cost per FIFO 6 = 34*6.6 7 = $ 224.4 9 Cost of

23 Average Cost 24 Cost of ending inventory 25 = No. of units in ending inventory* average cost 26 = 34*(1395.9/214) 27 = $ 2